OpenStudy (anonymous):

hi can you help me with this question grandpa wants to share his money among his grandsons if he gives them 30,000 each one he will be left with 60000, if he gives them 50,000 he will need an additional 100,000. how much money grandpa wishes to share among his grandsons?

OpenStudy (hoblos):

let's suppose first that the total money is y and the number of grandsons is x so we need to translate the given to a system of 2 equations with 2 unknowns x & y

OpenStudy (anonymous):

Is it how much money he wants to spend if he gives them 30,000 each or if he gives them 50,000 each?

OpenStudy (hoblos):

"if he gives them 30,000 each one he will be left with 60000" this means that if he gives each of the x grandsons 30000 , he would pay y-60000 so 30000x = y - 60000 this is the first equation can you fibd the secod one ?

OpenStudy (anonymous):

hey hoblos i see

OpenStudy (anonymous):

But the grandpa has 8 grandsons. Because if he gives them each 30,000, he will have 60,000 left and once he gives 50,000 each, he needs 100,000 more. So, a 20,000 increase per person means that there is an increase of 160,000. And 160,000/20,000 = 8.

OpenStudy (anonymous):

i dont get why that ecuation has both incognites on it

OpenStudy (anonymous):

hey so it woulb be something like this 50,000x = y + 100,000

OpenStudy (anonymous):

for the second ecuation

OpenStudy (hoblos):

yeah you are right, now you have to solve the system

OpenStudy (anonymous):

ok so this is the system

OpenStudy (anonymous):

30,000x = y - 60,000 50,000x = y + 100,000

OpenStudy (hoblos):

yep this is it, do you know how to solve it ?

OpenStudy (anonymous):

i dont quite remember.

OpenStudy (hoblos):

solve one of the 2 equations for y, then replace it in the other and solve for x

OpenStudy (anonymous):

so i have to rearrange the equations like this y=30,000x + 60,000 y=50,000x - 100,000

OpenStudy (hoblos):

yeah now set y=y and solve for x

OpenStudy (anonymous):

so 30,000x + 60,000 = 50,000x - 100,000

OpenStudy (hoblos):

correct

OpenStudy (anonymous):

30,000x - 50,000x = -60,000-100,000 20,000x = -160,000 x= -80,000 i got this but im getting a negative value is that right

OpenStudy (hoblos):



OpenStudy (hoblos):

there must be a "-" there

OpenStudy (anonymous):

oh my friend thanks for pointing that out :)

OpenStudy (hoblos):

so we get 20,000x = -160,000 now divide both sides by 20,000

OpenStudy (anonymous):

so yeah answer is 300,000 thanks my friend.

OpenStudy (hoblos):

yeah your answer is correct good work

OpenStudy (anonymous):

i really helped me knowing which were the things i was being asked money=y and grandsons=x

OpenStudy (anonymous):

Thank you so much.

OpenStudy (hoblos):

glad to help! :)